### v1

# Interpretability and Generalization Bounds for Learning Spatial Physics

Abstract
While there are many applications of ML to scientific problems that look promising, visuals can be deceiving. For scientific applications, actual quantitative accuracy is crucial. This work applies the rigor of numerical analysis for differential equations to machine learning by specifically quantifying the accuracy of applying different ML techniques to the elementary 1D Poisson differential equation. Beyond the quantity and discretization of data, we identify that the function space of the data is critical to the generalization of the model. We prove generalization bounds and convergence rates under finite data discretizations and restricted training data subspaces by analyzing the training dynamics and deriving optimal parameters for both a white-box differential equation discovery method and a black-box linear model. The analytically derived generalization bounds are replicated empirically. Similar lack of generalization is empirically demonstrated for deep linear models, shallow neural networks, and physics-specific DeepONets and Neural Operators. We theoretically and empirically demonstrate that generalization to the true physical equation is not guaranteed in each explored case. Surprisingly, we find that different classes of models can exhibit opposing generalization behaviors. Based on our theoretical analysis, we also demonstrate a new mechanistic interpretability lens on scientific models whereby Green’s function representations can be extracted from the weights of black-box models. Our results inform a new cross-validation technique for measuring generalization in physical systems. We propose applying it to the Poisson equation as an evaluation benchmark of future methods.
1 Introduction
The development of robust and rigorous a priori estimates for many numerical methods is a major victory of the past century. These estimates ensure geometric flexibility and convergence guarantees, laying the groundwork for complex physics simulations across a wide range of applications (Ciarlet, 2002; LeVeque, 2007). The foundation of all these results is the humble 1D Poisson equation, studied by virtually all numerical analysts and serving as the theoretical bedrock for computational physics.
Data-driven machine learning has risen dramatically recently as a new way to model previously intractable systems (Raissi et al., 2019; Lu et al., 2019). While ML/SciML is already successfully deployed in disparate domains from autonomous robotics to climate simulations, its theoretical foundations are crucial. Scientific applications are unique in the need for numerical precision and accuracy: even small percentage errors in parameters can be impactful. Recent works have advanced the theoretical understanding learning time-dependent differential equations (Queiruga et al., 2020; Ott et al., 2021; Zhu et al., 2022; Krishnapriyan et al., 2023), but thus far no comprehensive theory exists for learning the spatial component. Just as a foundational textbook chapter on classical numerical analysis rigorously studies solving the Poisson equation, this paper rigorously studies learning the Poisson equation to build the same theoretical foundations for ML. We combine traditional theoretical and empirical analysis with modern lenses, such as mechanistic interpretability through direct parameter inspection.
Figure 1: Expectations when learning a black box linear model, ${\bm{u}}={\bm{W}}{\bm{f}}$. Despite some visual similarity to the Green’s function operator (top), naive approaches are not guaranteed to converge to the true operator. The issue lies in the sampling procedure of the training data. The naive approach of using polynomial analytical solutions (illustrated as cubic polynomials) can achieve machine precision MSE loss, but even with infinitely many examples, will not converge to the general solution (center). It is possible to alter the construction of the training data, i.e. by using random piecewise linear functions solved via the finite element method (FEM) (bottom) to recover the true operator. This further allows for the extraction of a familiar discrete differential equation operator (right).
The canonical 1D Poisson equation on the unit interval with homogeneous Dirichlet boundary conditions is given by
$-k\frac{d^{2}u}{dx^{2}}=f(x),\qquad u(0)=u(1)=0.$ (1)
where $u(x)$ the solution field, $f(x)$ the forcing function, and $k$ the material property, e.g. thermal conductivity. The solution to Eq. (1) can be obtained by Green’s function which maps a function $f(x)$ to solution $u(x)$ as follows:
$u(x)=\int_{0}^{1}f(s)G(s,x)\,\mathrm{d}s+(u_{1}-u_{0})x+u_{0}$ (2)
The Green’s function to the Poisson differential equation (DE) is:
$G(s,x)=\begin{cases}(1-s)x&\text{if}\quad x<s\\ (1-x)s&\text{if}\quad x\geq s\end{cases}.$ (3)
In the ML problem, the differential equation is typically unknown and we are given discrete data of forcing functions $f(x_{i})$ and solutions $u(x_{i})$. We seek to train a model to approximate a solution operator that allows us compute unknown $u$ given unseen $f$. The domain of this problem is continuous, but measuring the fields introduces a discretization: values of $u(x_{i})$ and $f(x_{i})$ corresponding to measurements of the fields on a grid $x_{i}$. In practice, this discretization arises from a sensor grid or a camera, such that each grid point is analogous to a pixel in an image. For simplicity, we only consider uniform spacing of $\Delta x$. The simplest black-box model is a linear solution operator given by
${\bm{u}}={\bm{W}}{\bm{f}}$ (4)
where ${\bm{W}}$ is of shape $N_{\text{grid}}\times N_{\text{grid}}$ and can be trained via the usual gradient descent procedure on a mean square error loss, $\mathcal{L}=\sum\|{\bm{u}}-{\bm{W}}{\bm{f}}\|^{2}_{2}$.
A natural question arises as to what sort of matrix this model learns. We illustrate the impetus for this paper in Fig. 1. At first, we constructed a dataset using analytical solutions with cubic forcing functions, and achieved low train and test error on that space. The learned matrix (second row) qualitatively resembles the analytically derived matrix from Green’s function (first row). On further inspection, the learned matrix is quite far from the “truth”. The key insight is to take extra care on how the dataset is constructed. In general, the model will appear uninterpretable and fail to generalize to out-of-distribution (OOD) eval datasets. However, with careful construction of the dataset (bottom of Fig  1), we can actually recover both a full Green’s function discretization, ${\bm{A}}$ and a discretization of the differential operator, ${\bm{L}}={\bm{A}}^{-1}$.
To formalize, we define generalizability as achieving a low error, $\mathcal{L}()$, when sampling on a different the distribution than the training data. After minimizing $\mathcal{L}(W,\{u,f\}\sim\mathcal{F}_{train})$, we then desire:
$\mathcal{L}(W,\{u,f\}\sim\mathcal{F}_{\text{test}})\lesssim\mathcal{L}(W,\{u,f\}\sim\mathcal{F}_{\text{train}})$ (5)
when $\mathcal{F}_{\text{test}}\neq\mathcal{F}_{\text{train}}$. The usage of $\lesssim$ denotes a small “wiggle room” tolerance heuristic. We observe that the OOD error can be orders of magnitude larger than the observed training error, $\mathcal{L}_{\text{test}}\gg 10^{8}\times\mathcal{L}_{\text{train}}$.
We design a new cross-validation procedure that trains a given ML model with a dataset on a subspace, and then evaluates the error on datasets generated on different subspaces. We construct datasets by varying the function class of solutions $f\sim\mathcal{F}(type,p)$ and the grid spacing $\Delta x$ of the discretization. We consider four different modeling approaches make the following contributions:
  1. 1.
    Underparameterized models with an inductive bias of known physical equations. We demonstrate both empirically and theoretically that the estimated $w$ parameter converges to the true parameter $k$ at a rate dependent upon the grid spacing $\Delta x$, finite difference order $q$, and which increases with the order of training function basis $p$.
  2. 2.
    Exact parametrization for linear problems. We derive a true solution ${\bm{A}}$, and theoretically and empirically demonstrate that the learned weights only converges to a subspace of the operator based on the basis of the training data. The convergence is not bound by $\Delta x$.
  3. 3.
    Overparameterized neural models. We empirically demonstrate that deep models are not guaranteed to generalize outside of the training data. Subspace generalization is also not guaranteed. It is possible to extract Green’s function operators from black-box models.
  4. 4.
    Methods designed for differential equation learning. We observe that the DeepONet and Fourier Neural Operator also do not generalize well and behave qualitatively similarly to deep linear models in our experiments.
Counterintuitively, we discover that the data requirements and generalization results for each class of model are different and even opposite in some cases.
2 Background
Applications of ML in science relevant to discovering unknown dynamics from data generally span across “white-box” to “black-box”. White-box methods, such as SINDy and symbolic regression, allow for the obtainment of closed-form, analytical formulas (Rudy et al., 2017; Udrescu & Tegmark, 2020). Black-box methods such as Neural Operators or DeepONets generally use NNs to learn dynamics (Raissi et al., 2019; Lu et al., 2019), exchanging interpretability for flexibility. Between these two approaches, techniques have been developed that respect physical constraints (Hansen et al., 2023b) such as conservation laws (Hansen et al., 2023a; Patel et al., 2022) or other fundamental identities (Trask et al., 2022). Extending even beyond these, “world models” seek to replace simulation environments (Ha & Schmidhuber, 2018) with ML models operating directly on raw state-action observations data. Massive video-action transformer models are now commonly used in deployed self-driving car systems (Goff et al., 2025) and are expanding to other domains such as surgical robotics (Schmidgall et al., 2024). These applications require high quantitative accuracy and generalization. Even a 1% error on a latent physical parameter can have a grave impact.
Interpretability of internal model mechanics has become a subject of intense interest as ML models become more capable and general purpose. The phase change in deep models between memorization to generalization has been identified as the grokking phenomenon, with similar works detailing of watching weight matrix evolution toward prescribed algorithms (Power et al., 2022; Nanda et al., 2023). In a similar vein as scientific interpretability, transformers have been shown to be able to express and converge to algorithms for linear regression Zhang et al. (2024); Liang et al. (2024).
Recent strides in numerical analysis of ML have demonstrated that Neural ODEs and residual connections can overfit to time series data and may not learn true continuous models (Queiruga et al., 2020; Ott et al., 2021; Sander et al., 2022). Furthermore, it has been proven that higher order time discretization can elicit generalization(Zhu et al., 2022; Krishnapriyan et al., 2023). For spatiotemporal models, Krishnapriyan et al. (2021) demonstrated that PINNs will fail without controlled training strategies. The usage of Green’s functions in machine learning has been examined and used to prove data requirements for elliptic equations (Gin et al., 2021; Boullé et al., 2022, 2023). Other interpretations of SciML mechanisms as computational mathematics techniques include drawing analogies between transformers and quadrature, nonlocal operators, or kernel functions (Cao, 2021; Nguyen et al., 2022; Yu et al., 2024).
3 Theoretical Results
3.1 Dataset construction
To explore how the origin of training data impacts the model, we construct different datasets by sampling from different function spaces, assuming no noise. Let $\mathcal{F}(type,p)$ denote a function space of a type (polynomial, FEM, cosine, sine, etc.) with $p$ terms. Given the function class, the dataset is then constructed by sampling functions and evaluating on a grid with points $x_{i}\in i\Delta x$ with $\Delta x=1/N_{\text{grid}}$ (e.g. the spacing the sensor grid). Entries in the dataset are $\{x_{i},f(x_{i}),u(x_{i})\}$ from functions sampled from $f\sim\mathcal{F}(type,p)$.
In the theoretical analysis, we consider functions constructed from basis sets. We implement four distinct function classes for experiments with variable $p$. The first class is polynomials, $f(x)=\sum_{i=1}^{p}c_{i}x^{i}$. The Cosine and Sine datasets are defined by $f(x)=\sum_{i=1}^{p}c_{i}\cos(i\pi x)$ and $f(x)=\sum_{i=1}^{p}c_{i}\sin(i\pi x)$ respectively. For these classes, functions $f$ are drawn by sampling coefficients from a distribution $c_{i}\sim U[-1,1]$, and then $u(x)$ is analytically solved. The final function class, “FEM”, is constructed from piecewise linear $f$ with sampled node values $f_{i}\sim\mathcal{N}(0,1)$. The finite element method (FEM) provides analytically exact solutions for $u(x_{i})$ at node points in 1D. In this dataset, each point value ${\bm{f}}_{i}$ is independently random. The corresponding solution $u$ is determined from this specific $f$ (Fig. 1), which is distinct from adding noise to the observations of $f$ and $u$.
3.2 Parameter learning with a numerically discretized differential equations
Now consider the case where one has identified the form of the DE, through either prior knowledge or a system identification method, but do not know its parameters. The task is to fit a discretization of the Poisson equation for the single scalar parameter $w$, whose true counter part is $k$ in (1), to satisfy the equation $-w\mathrm{d}^{2}u(x_{i})/dx^{2}\approx f(x_{i})$ over the data. The DE is approximated using a finite difference stencil with the given $\Delta x$. E.g., the three-point stencil, referred to as FD-2, is
$\frac{d^{2}u}{dx^{2}}=\frac{u_{i-1}-2u_{i}+u_{i+1}}{\Delta x^{2}}+\mathcal{O}(\Delta x^{2}).$ (6)
It approximates the derivative with a local truncation error proportional to the grid size at a rate $q=2$. We can estimate $w$ given the data $\left\{u,f\right\}$ by analytically solving the minimization problem of the mean square error (MSE) loss over $N$ examples
$w=\arg\min_{w}\frac{1}{N}\sum_{n=1}^{N}\sum_{i=0}^{N_{grid}}\left(\frac{w}{\Delta x^{2}}\left(u_{i-1}^{(n)}-2u_{i}^{(n)}+u_{i+1}^{(n)}\right)+f_{i}^{(n)}\right)^{2}$ (7)
where superscript $(n)$ indexes a training example. The following theorem shows that the relative error $e=|w-k|/|k|$ can be bounded under the assumption of data sampled from polynomials:
Theorem 1.
Learning the parameter $k$ using a finite difference stencil of order $q$ given polynomial training data of order $p$ on a grid of spacing $\Delta x$ results in $w=k$ when $p<q$ and an error
$\frac{|w-k|}{|k|}=\mu_{q}\Delta x^{q}+\sum_{m=q+1}^{p}\mu_{m}\Delta x^{m}\quad when\quad p\geq q$ (8)
for constants $\mu_{m}$ depending on the truncation error coefficients of the finite difference stencils.
The proof is in Appendix A.1. Even in the absence of numerical error, there is still an error between the learned parameter $w$ and true parameter $k$ based on the grid size $\Delta x$ and polynomial order of the dataset. Increasing the grid size $\Delta x$ increases the error proportional to $\Delta x^{q}$. Incrementing the polynomial order of the training data from $p-1$ to $p$ increases the error proportional to $\Delta x^{p}$.
3.3 Training Dynamics for a Linear Model
We wish to learn a matrix ${\bm{W}}$ which converges to the action of a solution operator ${\bm{A}}$. Recall the Green’s function (2) for the Poisson problem, and its discretized problem (4). It is necessary to also discretize the Green’s function relative to the discrete measurement space. The continuous function can be constructed by $f(x)=\sum_{i}{\bm{f}}_{i}\phi_{i}(x)$ where $\psi_{i}$ is a basis of the grid discretization. Basis functions $\psi_{i}(x)$ could be piecewise constant representing camera pixels, or piecewise linear, representing interpolation between point sensors. The discrete solution operator is thus,
[…]
We let ${\bm{A}}_{ij}$ be the matrix in the brackets such that ${\bm{u}}={\bm{A}}{\bm{f}}$. The MSE loss over the dataset for this model is
$\mathcal{L}=\frac{1}{2N}\sum_{n=1}^{N}\left\|{\bm{W}}{\bm{f}}^{(n)}-{\bm{u}}^{(n)}\right\|^{2}_{2}$ (10)
We consider the case where our data is only sampled uniformly over subspaces of $\mathbb{R}^{n}$, such that ${\bm{B}}$ is the matrix of rank $p+1$ of coefficients satisfying $f(x_{i})={\bm{f}}_{i}={\bm{B}}{\bm{c}}$. For example, a polynomial function space can use the Vandermonde matrix ${\bm{B}}_{ij}=[1,x_{i},x_{i}^{2}...x_{i}^{p}]$. By Eq. (9), there exists a matrix ${\bm{A}}$ such that all training and test data satisfies ${\bm{u}}={\bm{A}}{\bm{f}}$. The following result shows that out of distribution generalizations cannot be expected:
Theorem 2.
Applying gradient descent a training dataset $\{{\bm{u}}^{(n)},{\bm{f}}^{(n)}\}$ sampled from a function space $\mathcal{F}_{\text{train}}$, the model weights ${\bm{W}}$ will converge to projection of the true operator ${\bm{A}}$ onto the subspace of $\mathcal{F}_{\text{train}}$. Let the forcing functions be sampled by ${\bm{f}}^{(n)}={\bm{B}}{\bm{c}}^{(n)}$ where each component of ${\bm{c}}$ is sampled independently with $\mathbb{E}[{\bm{c}}_{i}]=0$. Optimizing the loss (10) using gradient descent with an initial condition ${\bm{W}}^{0}$ will converge to
${\bm{W}}^{*}={\bm{A}}{\bm{U}}{\bm{U}}^{T}+{\bm{W}}^{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T}).$ (11)
where ${\bm{U}}$ is the left orthogonal basis of ${\bm{B}}$ of shape $N_{\text{grid}}\times(p+1)$.
The proof is in Appendix A.3. The learned matrix will converge to the projection of the true operator ${\bm{A}}$ onto the subspace of data $F$, ${\bm{A}}{\bm{U}}{\bm{U}}^{T}$, plus an error/noise that is orthogonal to the subspace of the data relative to the initial weights ${\bm{I}}-{\bm{U}}{\bm{U}}^{T}$. The latter “noisy” aspect could be mitigated by regularization such as weight decay, converging to ${\bm{W}}={\bm{A}}{\bm{U}}{\bm{U}}^{T}$ but not ${\bm{A}}$ exactly.
The above result is irrespective of the amount of data and the discretization, and as such, is actually a quite pessimistic result. In particular, ${\bm{A}}$ is only learned if and only if the dimension of subspace of ${\bm{f}}$ is of equal rank with ${\bm{A}}$. The finer the discretization, the greater the “quality” of the data needed (e.g. dimension of the subspace); infinite quantity of data is not sufficient.
4 Experiments
We examine six different types of ML models. We validate our theoretical results by comparing analytical expressions with experimental results. We then demonstrate that generalization can be inferred from data with careful cross-validation. Finally, we demonstrate the interpretation of model weights as the Green’s function operator. App. B details software, hardware, and hyperparameters. App.  D contains additional results on grid size.
To construct the datasets, we use the spaces described in Section 3.1. There are 25 different function classes considered: one piecewise linear dataset (FEM) and $p\in[1,8]$ for polynomial, sine, and cosine spaces. A model is trained on one dataset, and then evaluated on the other 24 datasets with the same $N_{grid}$. For each choice $(type,p,N_{\text{grid}})$, we generate 10,000 examples to represent the infinite data limit. Each dataset is normalized to have unit norm in $u$ to (1) minimize the effects of numerical ill-conditioning and round-off errors in the $p\geq 5$ datasets with extreme magnitudes, and (2) to make the mean squared errors (MSE) comparable across datasets.
4.1 Comparison of Empirical Errors to Analytical Theory
Figure 2: Comparison of numerical experiments to analytically derived ML parameters from Theorems 1 and 2 trained on polynomial spaces with $N_{grid}=22$. (Left) The observed error for the linear matrix matches predictions for lower order $p$. (Middle and right) For the finite difference model, the analytical assumptions overestimate the error, but the trend matches with observations.
The theoretical results can be used to compute predicted optimal parameters, allowing for a priori estimations of errors as shown in Fig. 2. The equation from the proof of Theorem 1 can analytically solve expressions of $w$ as in App. A.2. The equation from Theorem 2 can be used to analytically predict ${\bm{W}}(p)$ by symbolically computing ${\bm{A}}$ and ${\bm{U}}$ using the procedure in A.5. The observed qualitative behavior matches that predicted by the theorems.
4.2 OOD Dataset Generalization
(a) Finite Different: $w[\mathtt{FD2}({\bm{u}},\Delta x^{2})]={\bm{f}}$
(b) Linear: ${\bm{u}}={\bm{W}}{\bm{f}}$
(c) Deep Linear: ${\bm{u}}={\bm{W}}_{2}({\bm{W}}_{1}{\bm{f}}+{\bm{b}}_{1})+{\bm{b}}_{2}$
(d) MLP: ${\bm{u}}={\bm{W}}_{2}\sigma({\bm{W}}_{1}{\bm{f}}+{\bm{b}}_{1})+{\bm{b}}_{2}$
Figure 3: Heatmap of evaluation error for different training and testing combinations across four model types. Each row represents one model trained the dataset labeled on the y axis, and each column shows the MSE on corrosending the test set. The dashed lines separate the function class families into blocks. Within the each block, the cells to the left of a cell are function subspaces. The color map represents the MSE on an entire test dataset on a log scale: sharp transitions in color values correspond to ×$\times 10^{10}$ increase in MSE. Note the blue lower triangular blocks in 3(b) and 3(c): these are cases where the model generalizes satisfying $\mathcal{L}(w,\mathcal{F}_{test})\lesssim\mathcal{L}(w,\mathcal{F}_{train})$ when $\mathcal{F}_{\text{test}}\subseteq\mathcal{F}_{\text{train}}$.
We conduct the training procedure for six types of models on datasets constructed with $N_{\text{grid}}=22$. We train each type of model from scratch using each of the 25 different datasets as the training dataset. After training, each model is then evaluated on the 24 other datasets that unseen during its training. This results in a $25\times 25$ grid of MSEs for each model, shown in Fig. 3 and 4 as heatmaps. The model with the best training MSE from 5 runs is chosen to its corresponding row in the heatamp. The implementation details and hyperparameters for all experiments are provided in App. B. Line plots of the same data with error bars over multiple seeds are included in App. C.
Finite Difference Model
As predicted by Theorem 1, the test error is observed in Fig. 3(a) to increase as the polynomial order of the training data increases. The error is confounded by both the error from the finite difference and the error arising from the $|w-k|$ term. The error increases by both generalizing across higher order functions (the former term), moving to the right in the heatmap, and training on higher order functions (the latter term), moving down in the heatmap. Fig. 2 and Fig. 14 in App. D further explore this error in more settings. The error for the FEM dataset is orders of magnitude higher as it does not satisfy smoothness requirements for finite differences.
Linear Model
We turn to Theorem 2 to empirically verify its predictions of ${\bm{W}}$ in Fig. 3(b). We initialize $W_{0}=0$, and do not use regularization. For each training run, the test error rises sharply when the test data are no longer in a subspace of the training data. There are three clear lower-triangular blocks in the heatmap representing the three families of nested subspaces, providing evidence for generalization across many types of function subspaces. Increasing polynomial orders decrease the MSE on the Cosine and Sine datasets (forming pyramid structures in the top middle and right.) The piecewise linear function space along the top row contains the $p=1$ space exactly, and consistently yields low MSE in all models.
Deep Linear Models and Shallow Neural Networks
We apply nonlinear parameterizations to the linear problem by training two different deep learning models, a deep linear model and a shallow neural network. Deep linear models are a tractable stepping stone between linear models and MLP (Saxe et al., 2013; Arora et al., 2018) to understand nonlinear training dynamics. Both models are trained with a larger hidden dimension than the input and output size (App. B). As seen in Figs. 3(c) and 3(d), both models do not exhibit generalization behavior consistently. Both achieve low test error on $p_{\text{train}}$, but the test error sharply increases for any other dataset where $p_{\text{test}}\neq p_{\text{train}}$. The MLP is strongly diagonal, showing that the model will overfit to the training data, and not generalize in any case. The Deep Linear model shows mixed behavior: it exhibits subspace generalization on the sine and cosine datasets but not on the polynomial datasets. As with the linear model, some cross-subspace generalization is observed when the models are trained with piecewise linear functions.
DeepONet and Neural Operators
(a) DeepONet: $u(x)=trunk({\bm{f}})\cdot branch(x)$
(b) Neural Op.: $u(x)=\mathtt{fft}^{-1}(R_{\phi}\cdot\mathtt{fft}({\bm{f}}))$
Figure 4: Generalization to out-of-distribution test datasets for the DeepONet and Fourier Neural Operator. The generalization patterns are similar to the Deep Linear models in Fig. 3.
We now investigate two models that are designed for learning PDEs: the DeepONet Lu et al. (2021a) and the Fourier Neural Operator Kovachki et al. (2021). The results are displayed in Fig. 4. The DeepONet and the Neural Operator support flexible $x$-coordinate distributions, but for both of these methods we still consider the regular grid, and do not consider grid adaptation of these methods. The DeepONet shows block lower-triangular structure similar to that of the linear model. The diagonal bands (evaluation on the training distribution) exhibit lower error values, showing a slight overfitting that does not strictly satisfy the subspace generalization. The dips in the trend are clear in Fig. 11 in App. C, where the model could satisfy $\mathcal{L}_{test}\lesssim\mathcal{L}_{train}$ tolerance depending on the application. The Neural Operator has a similar generalization pattern, but did not achieve a low training MSE on all higher order function classes.
4.3 Extracting Green’s functions from black-box models.
Figure 5: Visualization of the parameter matrix ${\bm{W}}$ of the linear model as the dataset order increases. The top row displays one row $W_{i1}$ of the matrix, the row vector corresponding to the impulse at point $x=\Delta x$. It can be seen to converge to the shape of a Green’s function. The middle row shows the entire matrix as a heatmap, and the bottom displays its inverse. The learned matrix converges incrementally towards the expected matrix (visualized from left to right) as terms are incrementally added to the training dataset. When the model is trained on the FEM dataset, it is possible to inverting the learned ${\bm{W}}$ and yield a banded structure of a local stencil. Parameters learned on all other dataset types cannot be inverted reliably.
Figure 6: Visualization of test function evaluations, $\text{Model}({\bm{e}}_{i})$, on black-box models can reveal Green’s function structure. The top row shows one test function evaluation at the node $x=\Delta x$ (each column has a different y-scale), and the bottom row shows a matrix constructed from model evaluations for all test functions. The colormap limits on the bottom heatmaps are set to match the y-scale of the corresponding plot above them.
The Green’s function is defined as the impulse response of the solution operator. Given a trained model, the action of the Green’s function can be approximated by applying a test function $\hat{f}_{k}=\delta_{kj}={\bm{e}}_{j}$ (the one-hot vector at index $j$) to extract the structure using the relation
${\bm{A}}_{ij}\leftrightarrow Model({\bm{f}}={\bm{e}}_{j})_{i}.$ (12)
For the linear model, this is the same as inspecting the weight matrix, ${\bm{A}}_{ij}\leftrightarrow{\bm{W}}_{ij}$. Fig. 5 elaborates on Fig. 1 to show the convergence of the weight matrix to the derived analytical solution for increasing modes (or complexity) of data. For lower orders $p$, the extracted functions are inscrutable, even though the model achieves low error on the training set. Training the linear model on the FEM dataset is sufficient to recover an accurate Green’s operator ${\bm{A}}$. It is even possible to invert the weight matrix (${\bm{W}}^{-1}=\hat{{\bm{L}}}$) to uncover a local stencil structure, which is evident in its qualitatively tridiagonal appearance.
Fig. 6 shows that it is also possible to extract the Green’s functions for deep models by evaluating the nonlinear models on one-hot inputs ${\bm{f}}=\textbf{e}_{j}$. For $p=3$, there is no obvious evidence of the expected structure, even the model achieves low MSE on the training set. The expected structure is extracted using the MLP and DeepONet when they are trained on FEM data. The structure is somewhat evident for the Fourier Neural Operator with the $cos[7]$ training run, but no others.
5 Conclusion
Learning spatially dependent problems is limited by the underlying function space of the training data and the observation discretization. We theoretically proved that this barrier exists on the simplest problem — the Poisson problem — with the simplest models, parameter fitting finite differences and linear models. We illustrated that carefully choosing the training data can mitigate this issue and that it is possible to generalize to other function spaces, which can inform data collection and cross-validation techniques. As another form of validation, black-box models can be directly interpreted as integrated Green’s functions when they generalize effectively across many test domains. However, this is still pessimistic and presents a headwind for discovering unknown physics from real world data, given the inherent limitations in real world data collection.
We demonstrated this behavior across a wide range of representative model types, from simple parameter fits to models designed for PDE learning, but it may be possible to design a modeling technique or data augmentation strategy to mitigate this issue. We did not consider any methods that incorporate a PDE loss or physics-informed loss, such as Physics Informed Neural Networks (PINNs) or Physics Informed Neural Operators (PINOs), because we were specifically studying the case of discovering the PDE operator from data without any prior knowledge. Our analysis of the finite different fitting suggest that even the strongest physical priors may not mitigate the impact of the training data distribution. Even more confoundingly, different types ML models can exhibit different and sometimes opposing generalization behaviors.
While our theory required a priori knowledge of how the dataset was constructed, our results suggest that extracting generalized scientific knowledge may require new methods capable of compensating for these limitations by discovering the underlying subspace of the data. We propose this methodology to be an evaluation benchmark on future development of learning physics. As we have seen that even the Poisson equation can pose a challenge, we recommend that future proposals for PDE learning techniques measure cross-set generalization on Poisson data.
Limitations
This work is an initial theoretical exposition of the newly discovered behavior, and leaves many questions open. We only performed the exploration and proved theorems on a single differential equation to focus on demonstrating the result across different models. Future work should extend this analysis to more complex physical systems and ML models, and construct a similar evaluation on real-world data. We do not currently have any recommendations for how to solve the lack of generalization; we leave this challenge open to future research to utilize the proposed evaluation method to improve existing techniques and innovate for learning of physical systems.
Appendix A Proofs
A.1 Proof of Theorem 1
Proof.
The training forcing function and corresponding analytical solution is given by
$f^{(n)}(x)=-k\sum_{m=0}^{p}c_{m}x^{m},\quad u^{(n)}(x)=\sum_{m=-2}^{p}\frac{c_{m}x^{m+2}}{(m+2)(m+1)}$ (13)
for coefficients $c_{m}$. Recall that a finite difference stencil (LeVeque, 2007) of accuracy order $q$ will satisfy
$FD_{q}(u,\Delta x)=u^{\prime\prime}+\mathcal{O}(\Delta x^{q}).$ (14)
Thus for a polynomial whose true derivative is $\sum c_{m}x^{m}$, the finite difference stencil can be written as a polynomial whose coefficients change starting at its accuracy order $q$,
$FD_{q}(u,\Delta x)=\sum_{m=0}^{q-1}c_{m}x^{m}+\sum_{m=q}^{p}c_{m}(1+\zeta_{m}\Delta x^{m})x^{m}:=\sum_{m}b_{m}x^{m}$ (15)
where $\zeta_{m}$ is a bound of the coefficient of error introduced for each term $m$ starting at $q$ by Taylor’s theorem. These error terms are are dependent upon the finite difference stencil; the following section illustrates two examples. This proof is for the general case. For the usual 3 point stencil, $q=2$. Simplifying the loss from (29)
[…]
where $b_{m}=c_{m}$ for $m<q$ and $b_{m}=c_{m}(1+\zeta_{m}\Delta x^{m})$ for $m\geq q$., Expanding the square, applying the integral and simplifying using the definition of $b_{m}$
[…]
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}(wb_{m}-kc_{m})(wb_{j}-kc_{j})\frac{1}{m+j+1}\right]$ (19)
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)c_{m}c_{j}\frac{1}{m+j+1}\right]$ (20)
$=\sum_{j=0}^{p}\sum_{m=0}^{p}(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)\frac{\mathbb{E}_{c}\left[c_{m}c_{j}\right]}{m+j+1}$ (21)
where we let $\xi_{m}:=\zeta_{m}\Delta x^{m}$. Since $c_{m}$ are iid, then $\mathbb{E}[c_{m}c_{j}]=\mathbb{E}[c_{m}^{2}]\delta_{mj}$. This lets us only consider the diagonal terms of the sum,
$\mathcal{L}$ $=\sum_{m=0}^{p}(w(1+\xi_{m})-k)^{2}\frac{\mathbb{E}_{c}\left[c_{m}^{2}\right]}{2m+1}.$ (22)
Taking the derivative with respect to $w$
$\frac{d\mathcal{L}}{dw}$ $=\sum_{m=0}^{p}2(1+\xi_{m})(w(1+\xi_{m})-k)\frac{\mathbb{E}_{c}\left[c_{m}^{2}\right]}{2m+1}.$ (23)
Setting $d\mathcal{L}/dw=0$, the minima is at
$w$ $=\frac{\sum_{m=0}^{p}(1+\xi_{m})\frac{\mathbb{E}_{c}\left[c_{m}^{2}\right]}{2m+1}}{\sum_{m=0}^{p}(1+\xi_{m})(1+\xi_{m})\frac{\mathbb{E}_{c}\left[c_{m}^{2}\right]}{2m+1}}k.$ (24)
By assumption that the expectations on the second moment are identical for each coefficient $c_{m}$, we can simplify and then break up the sum,
$w$ $=\frac{\sum_{m=0}^{p}(1+\xi_{m})\frac{1}{2m+1}}{\sum_{m=0}^{p}(1+\xi_{m})(1+\xi_{m})\frac{1}{2m+1}}k$ (25)
$=\frac{\sum_{m=0}^{q-1}\frac{1}{2m+1}+\sum_{m=q}^{p}(1+\zeta_{m}\Delta x^{m})\frac{1}{2m+1}}{\sum_{m=0}^{q-1}\frac{1}{2m+1}+\sum_{m=q}^{p}(1+2\zeta_{m}\Delta x^{m}+\zeta_{m}^{2}\Delta x^{2m})\frac{1}{2m+1}}k$ (26)
Now by algebraically using this to compute the error term, we have
$\frac{w-k}{k}=\frac{-\sum_{m=q}^{p}(\zeta_{m}\Delta x^{m}+\zeta_{m}^{2}\Delta x^{2m})\frac{1}{2m+1}}{\sum_{m=0}^{q-1}\frac{1}{2m+1}+\sum_{m=q}^{p}(1+2\zeta_{m}\Delta x^{m}+\zeta_{m}^{2}\Delta x^{2m})\frac{1}{2m+1}}.$ (27)
Note that $w=k$ if $p<q$ as the numerator vanishes. Conversely, when $p>q$, $w\to k$ as $\Delta x\rightarrow 0$ at a rate of $\Delta x^{q}$ by polynomial division. Furthermore, for a fixed $\Delta x$, increasing $p$ Each additional term adding $p$ adds additional terms: at the very least, the term $\xi_{m}\Delta x^{2m}$ is always positive, so the magnitude of that term will always increase, even if $\zeta_{m}\Delta x^{m}$ has favorable canceling out.
There, the dominant terms are the numerator,
$\frac{w-k}{k}=\sum_{m=q}^{p}(\mu_{m}\Delta x^{m})$ (28)
where the particular coefficients $\mu_{m}$ are finite constants that can be derived for particular finite difference rules. ∎
A.2 Concrete Example of Error for 3 point stencil
For a given stencil, the above procedure can be executed algebraically with fewer assumptions. As above, we consider the limit of infinitely many samples of functions $f,u$ with $c\sim C$ and integrating over $x\in[0,1]$.
$\mathcal{L}=\mathbb{E}_{c}\left[\int_{0}^{1}\left(wFD(u,\Delta x)-ku^{\prime\prime}\right)^{2}\mathrm{d}x\right].$ (29)
Thus, given a training set of $u$ and $f$ generated by sampling coefficients $c_{n}$, the inner value of (7) is
$\frac{w}{\Delta x^{2}}\left(\left(\sum_{n=0}^{p+2}c_{n}(x-\Delta x)^{n}\right)-2\left(\sum_{n=0}^{p+2}c_{n}x^{n}\right)+\left(\sum_{n=0}^{p+2}c_{n}(x+\Delta x)^{n}\right)\right)\\ -k\sum_{n=2}^{p+2}n(n-1)c_{n}x^{n-2}$ (30)
meaning the interior of the MSE loss
$\mathcal{L}=\mathbb{E}_{c}\left[\int_{0}^{1}\left(\frac{w}{\Delta x^{2}}\sum_{n=2}^{p+2}c_{n}\left((x-\Delta x)^{n}-2x^{n}+(x+\Delta x)^{n}\right)\\ -k\sum_{n=2}^{p+2}n(n-1)c_{n}x^{n-2}\right)^{2}\mathrm{dx}\right].$ (31)
For $p=0$ and $p=1$, simplifying the finite difference stencil results in an equation where $w=k$ within the squared error term.
But when we consider $p=2$ with quadratic $f$ and quartic $u$, there is not full cancellation:
[…]
$w=\frac{3k\left(105\Delta^{2}+223\right)}{245\Delta^{4}+630\Delta^{2}+669}$ (33)
which leads an error term
$\frac{w-k}{k}=\frac{\Delta^{2}\left(-245\Delta^{2}-315\right)}{245\Delta^{4}+630\Delta^{2}+669}$ (34)
which satisfies the general result.
A.3 Proof of Theorem 2
Proof.
By construction of our dataset and ansatz on ${\bm{A}}$, the $n$th sample is ${\bm{f}}^{(n)}={\bm{B}}{\bm{c}}^{(n)}$ and ${\bm{u}}^{(n)}={\bm{A}}{\bm{B}}{\bm{c}}^{(n)}$, thus the loss is
$\mathcal{L}=\frac{1}{2N}\sum_{n=1}^{N}||{\bm{A}}{\bm{B}}{\bm{c}}^{(n)}-{\bm{W}}{\bm{B}}{\bm{c}}^{(n)}||^{2}$ (35)
where ${\bm{A}},{\bm{W}}$ are of size $N_{\text{grid}}\times N_{\text{grid}}$, ${\bm{c}}$ is of size $p$ and ${\bm{B}}$ is size $N_{\text{grid}}\times p$.
Clearly ${\bm{W}}={\bm{A}}$ is a solution, but if ${\bm{B}}$ is lower rank than ${\bm{A}}$ and ${\bm{W}}$, there is a nullspace in the least squares equation. Gradient descent moves in the direction of the gradient of the loss
$\frac{\partial\mathcal{L}}{\partial{\bm{W}}}=\frac{1}{N}\sum_{n=1}^{N}({\bm{A}}-{\bm{W}})({\bm{B}}{\bm{c}}^{(n)}({\bm{c}}^{(n)})^{T}{\bm{B}}^{T}).$ (36)
With the assumption of iid ${\bm{c}}$, then $\mathbb{E}[c_{m}c_{k}]=\mathbb{E}[c_{m}^{2}]\delta_{mk}$. For any sampling and basis set, we can rescale ${\bm{B}}_{:,m}$ such that $\mathbb{E}[c_{m}^{2}]=1$ to simplify the following. By linearity of expectation, we have that the expected gradient is simply
$\frac{\partial\mathcal{L}}{\partial{\bm{W}}}=({\bm{A}}-{\bm{W}}){\bm{B}}{\bm{B}}^{T}.$ (37)
The minima will satisfy
$0=({\bm{A}}-{\bm{W}}){\bm{B}}{\bm{B}}^{T}$ (38)
via the dynamics
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta({\bm{A}}-{\bm{W}}_{t}){\bm{B}}{\bm{B}}^{T}.$ (39)
With initial condition ${\bm{W}}^{0}$, we will show that the sequence converges to
${\bm{W}}^{*}={\bm{A}}{\bm{U}}{\bm{U}}^{T}+{\bm{W}}^{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})$ (40)
where ${\bm{U}}$ is the left $N_{\text{grid}}\times(p+1)$ eigenvector matrix of the SVD decomposition ${\bm{B}}={\bm{U}}{\bm{D}}{\bm{V}}^{T}$. Define the error at each step $t$ as
${\bm{E}}_{t}={\bm{W}}_{t}-{\bm{W}}^{*}$ (41)
The error can be decomposed into two components. ${\bm{E}}^{||}$ which is in the row space of ${\bm{B}}{\bm{B}}^{T}$ and ${\bm{E}}^{\bot}$ which is in the null space $null({\bm{B}}{\bm{B}}^{T})=I-{\bm{U}}{\bm{U}}^{T}$. The matrix ${\bm{B}}{\bm{B}}^{T}={\bm{U}}{\bm{D}}{\bm{V}}{\bm{V}}^{T}{\bm{D}}{\bm{U}}^{T}={\bm{U}}{\bm{D}}^{2}{\bm{U}}^{T}$. Since the gradient update is always in the row space of ${\bm{B}}{\bm{B}}^{T}$ this implies that ${\bm{E}}^{\bot}$ is constant 0 as the following calculation shows
${\bm{E}}_{0}^{\bot}$ $={\bm{W}}_{0}^{\bot}-{\bm{W}}^{*\bot}$
$={\bm{W}}_{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})-({\bm{A}}{\bm{U}}{\bm{U}}^{T}-{\bm{W}}^{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T}))({\bm{I}}-{\bm{U}}{\bm{U}}^{T})$
$={\bm{W}}_{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})-{\bm{W}}_{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})=0$
where we used ${\bm{U}}{\bm{U}}^{T}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})=0$. Thus, only considering the component parallel to ${\bm{U}}{\bm{U}}^{T}$, we perform the algebra on the updates to the error,
${\bm{E}}^{||}_{t+1}$ $={\bm{W}}_{t+1}^{||}-{\bm{W}}^{*||}=({\bm{W}}_{t+1}-{\bm{W}}^{*}){\bm{U}}{\bm{U}}^{T}$
$=({\bm{W}}_{t+1}-{\bm{A}}{\bm{U}}{\bm{U}}^{T}-{\bm{W}}_{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})){\bm{U}}{\bm{U}}^{T}$
$={\bm{W}}_{t+1}{\bm{U}}{\bm{U}}^{T}-{\bm{A}}{\bm{U}}{\bm{U}}^{T}$
$=({\bm{W}}_{t}+\eta({\bm{A}}-{\bm{W}}_{t}){\bm{B}}{\bm{B}}^{T}){\bm{U}}{\bm{U}}^{T}-{\bm{A}}{\bm{U}}{\bm{U}}^{T}$
$={\bm{W}}_{t}{\bm{U}}({\bm{I}}-\eta{\bm{D}}^{2}){\bm{U}}^{T}-{\bm{A}}{\bm{U}}({\bm{I}}-\eta{\bm{D}}^{2}){\bm{U}}^{T}$
Assuming the singular values of ${\bm{B}}$ are identical,
$=(1-\eta D^{2})({\bm{W}}_{t}^{||}-{\bm{W}}^{*||})$
$=(1-\eta D^{2}){\bm{E}}^{||}_{t}.$
Thus, for any $\eta<1/\max(D^{2})$, the components parallel to ${\bm{U}}{\bm{U}}^{T}$ goes to zero ${\bm{E}}^{||}_{t}=(1-\eta)^{t}{\bm{E}}^{||}_{0}\to 0$. As both error terms converges to 0, we have shown ${\bm{W}}_{t}\to{\bm{W}}^{*}$ as desired. ∎
A.4 Corollary to Theorem 2: A Loose Error Bound
A loose error bound naturally follows from this:
Corollary 1.
Assuming that the basis set has $p$ entries, the error $\|{\bm{W}}^{*}-{\bm{A}}\|_{F}/\|{\bm{A}}\|_{F}$ is dependent on the size of the basis set,
$\frac{\|{\bm{W}}^{*}-{\bm{A}}\|_{F}}{\|{\bm{A}}\|_{F}}\leq\sqrt{N_{\text{grid}}-p}\left(\frac{\|{\bm{W}}^{0}\|_{F}}{\|{\bm{A}}\|_{F}}+1\right)$ (42)
Proof.
Note that, ${\bm{W}}^{*}-{\bm{A}}={\bm{W}}^{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})-{\bm{A}}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})=({\bm{W}}^{0}-{\bm{A}})({\bm{I}}-{\bm{U}}{\bm{U}}^{T})$. The Frobenious norm satisfies the sub-multiplicative property, meaning $\|({\bm{W}}^{0}-{\bm{A}})({\bm{I}}-{\bm{U}}{\bm{U}}^{T})\|_{F}\leq\|{\bm{W}}^{0}-{\bm{A}}\|_{F}\|{\bm{I}}-{\bm{U}}{\bm{U}}^{T}\|_{F}$. The matrix ${\bm{U}}{\bm{U}}^{T}$ is rank $p$. The complementary matrix has rank $N_{\text{grid}}-p$. The F-norm for a rank $r$ orthogonal matrix is $\sqrt{r}$ by definition. Using the additive distribution of the F-norm, $\|{\bm{W}}^{0}-{\bm{A}}\|=\|{\bm{W}}^{0}\|+\|{\bm{A}}\|$. Assembling these gives the stated result. ∎
A.5 Corollary to Theorem 1: A Tight Error Bound
If the Green’s function is known, the basis ${\bm{B}}$ and analytical solution ${\bm{A}}$ can be computed to give a much tighter estimate, as is shown in Fig.2. The procedure is as follows. Compute a matrix of basis functions evaluated at each point in the domain,
${\bm{B}}_{ij}:=[\phi_{0}(x_{i}),\phi_{1}(x_{i}),\ldots,\phi_{p}(x_{i})]\quad x_{i}=i\Delta x$ (43)
and compute the SVD of the basis matrix:
${\bm{U}},{\bm{D}},{\bm{V}}:=\mathtt{SVD}({\bm{B}})$ (44)
Generate the Green’s function matrix by symbolically integrating the Green’s function using a compact basis:
${\bm{A}}_{ij}=\int_{0}^{1}\psi_{i}(s)G(s,x_{j})\mathrm{d}x$ (45)
where $\psi_{i}$ is a piecewise linear or a piecewise constant basis function centered on $x_{i}$. Then, the error starting at any initial matrix ${\bm{W}}^{0}$ can be directly computed as
$\|{\bm{W}}^{*}-{\bm{A}}\|_{F}\approx\|{\bm{W}}^{0}-{\bm{A}}{\bm{U}}{\bm{D}}{\bm{V}}^{T}\|_{F}$ (46)
where the computation for ${\bm{A}}$ has error turns from numerical roundoff and the arbitrariness of the choice of discretization basis $\psi(x)$.
Appendix B Procedure Details
To initially rule out errors from software drivers or numerical precision, we replicated the linear models and deep models in many hardware and software variations: Tinygrad, JAX, and PyTorch, on CPU, NVIDIA 4070 GPU, and Apple Silicon GPU, and varied precision in float32 and float64 on CPU. The ultimate experiments ran in float32 on GPUs on a Macbook Pro M3 Max. A single training run for any of the experiments take between 30 seconds to a few minutes on consumer hardware. A full sweep can take 20 minutes. All replications with five seeds takes about 48 hours of compute time on a single commodity laptop. Code scripts to replicate the experiments is included in the supplemental material.
Tuning of hyperparameters of hidden size and learning rate schedule was guided by achieving below $10^{-14}$ train MSE on a $p\geq 5$ function space. The most important aspect in each model was removing weight decay and having a schedule that terminates at a learning rate of $10^{-5}-10^{-6}$.
Finite difference experiments were performed using pysindy (de Silva et al., 2020; Kaptanoglu et al., 2022), running on CPU in float64 precision. While this library is meant for dynamics, the SINDy techniques will work for non-temporal spatial problems by simply setting the time derivative equal to the forcing function, $\mathtt{x\_dot}[i]={\bm{f}}_{i}$.
The DeepONet is trained using the DeepXDE library (Lu et al., 2021b). The Neural Operator is trained using the neuralop library (Kossaifi et al., 2024).
The polynomial, sine, and cosine datasets were generated using SymPy by symbolically solving the differential equation for $u(x)$ for the form of $f(x)$. The polynomial data is generated using a root-form construction by $f(x)=\Pi^{p}_{i=1}(x-r_{i})$ for specified $p=[1,2,\ldots,8]$ with roots $r_{i}\sim U[-1,2]$. It is easier to control for ill-conditioning than with the monomial basis for high degree polynomials. The sine and cosine data are generated using the normal series in the text. The piecewise linear datasets were generated using the finite element method using FEniCS (Baratta et al., 2023).
Hyperparameters and training parameters for the models are given in Tables 1, 2, 3, 4, and 5.
Table 1: Hyperparameters for Linear model training (see Section 4.2)
Hyperparameter Values
Batch size Full batch training on 10k examples
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-1 to 1e-6
Epochs 2,000
Table 2: Hyperparameters for Deep Linear model training (see Section 4.2)
Hyperparameter Values
Layers 1
Hidden dimension 100
Batch size Full batch training on 10k examples
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-1 to 1e-6
Epochs 2,000
Table 3: Hyperparameters for MLP training (see Section 4.2)
Hyperparameter Values
Layers 1
Hidden dimension 1024
Activation function Leaky ReLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-2 to 1e-6
Epochs 5,000
Table 4: Hyperparameters for Deep ONet training (see Section 4.2)
Hyperparameter Values
Layers 1
Branch x-tower dimensions [1, 256, 256]
Trunk F-tower dimensions [$N_{grid}$, 256, 256]
Activation function ReLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Step from 1e-3 to 1e-4, 1e-5, 1e-6 every 5,000 steps
Training steps 20,000
Table 5: Hyperparameters for Fourier Neural Operator training (see Section 4.2)
Hyperparameter Values
Layers 1
N modes 64
Hidden dimension 32
Activation function GeLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-2 to 1e-6
Epochs 5,000
Appendix C Repeatability of test set generalization
Each run was repeated 5 times with different random seeds to verify the repeatability of the results. No seed results in a change in behavior for any for the models. The results are shown in Figures 7, 8, 9, 10, 11, and 12.
The finite difference model and linear model will always optimize to the same solution, where the linear model will have different noise in the null space. With the nonlinear models, it is possible that a lucky seed results in better generalization. However, we did not observe any. Repeated experiments with different seeds are shown in Figures 7, 8, 9, 10, 11, and 12. Each line on this plot corresponds to training on one function class identified in the legend, which corresponds to a row in the heat map from Figures 3 and 4. Only ten rows are shown for clarity. The error bars represent min/max over 5 runs, and the line is the average of the five. The repeated runs all show the same qualitative trends of generalization and overfitting in the same regimes for each of the six types of models. The repeated training runs are consistent with low variation in MSEs. The Neural Operator has the largest variability in MSEs between runs.
Figure 7: Generalization errors for the Finite Difference SINDy model with five different seeds for each training run. The finite difference optimization is invariant to the seed, so the min and max are within machine precision. The lower y limit is truncated to $10^{-12}$: training on Poly 1, Sine 1 and Cosine 1 achieves error $\approx 10^{-38}$.
Figure 8: Generalization errors for the linear model with five different seeds for each training run. With different random seeds, the min and max on the eval errors are within machine precision for the linear model.
Figure 9: Generalization errors for the deep linear model with five different seeds for each training run.
Figure 10: Generalization errors for the MLP with five different seeds for each training run.
Figure 11: Generalization errors for the DeepONet with five different seeds for each training run.
Figure 12: Generalization errors for the Neural Operator with five different seeds for each training run.
Appendix D Grid size generalization
Figure 13: Convergence of the weight matrix $W$ as the grid size increases for two polynomial orders and piecewise linear forcing functions. The matrix structure is invariant to the discretization, however, each parameter scales proportional to $\Delta x$, as implied by Eq. 9.
Figure 14: Experiments on SINDy equation discovery as the grid size (x-axis) and training data polynomial function order (line color in legend) is changed. The number of parameters and equation libraries stays fixed as the number of points in the domain grows. 14(a): training data MSE (observable by the scientist). 14(b) : difference from true parameter (not observable). After convergence, the error tends to increase; we believe this is accumulation of floating point error as the points per example increases and not a fundamental limitation. I.e., with infinite precision we expect the error to go to 0.
While with time-dependent systems it is possible to measure $\Delta t$ between frames, we do not have a proposal for how to adapt models when $\Delta x$ in different examples may be variable and unknown. We consider the situation where we train and evaluate models on datasets produced at different grid spacings. (What if we bought camera A instead of camera B?) Note that $\Delta x$ is a property of the training data: decreasing $\Delta x$ in practice would require finer measurements by, for example, buying a better a camera with higher resolution.
In Fig. 13, we observe that the shape of the matrix does not change as the grid size increases: the model does not learn anything different. The polynomial datasets have the same number of coefficients for each grid size. However, the rank of “fem” dataset does increase as the piecewise linear functions are sampled the new grid.
For the finite difference method in Fig. 14, we do observe a difference in the error. As predicted by Theorem 2, the error on the learned parameter is proportional to both the grid size and polynomial order of the dataset. The figures also make it clear the order of the stencil does change the rate of convergence towards the true parameter. Increasing the order of accuracy of the finite difference approximation used is possible in practice, however, that increases the sensitivity to noise.